### v2

# A fully Bayesian approach for the imputation and analysis of derived outcome variables with missingness

Abstract
Derived variables are variables that are constructed from one or more source variables through established mathematical operations or algorithms. For example, body mass index (BMI) is a derived variable constructed from two source variables: weight and height. When using a derived variable as the outcome in a statistical model, complications arise when some of the source variables have missing values. In this paper, we propose how one can define a single fully Bayesian model to simultaneously impute missing values and sample from the posterior. We compare our proposed method with alternative approaches that rely on multiple imputation with examples including an analysis to estimate the risk of microcephaly (a derived variable based on sex, gestational age and head circumference at birth) in newborns exposed to the ZIKA virus.
Keywords: multiple imputation, missingness, Bayesian inference.
1 Introduction
‘Derived variables’ are variables that are constructed from one or more measured variables through established mathematical operations or algorithms. Examples of commonly used derived variables include body mass index (BMI), constructed from weight and height, and the 36-Item Short Form Health Survey Physical Component Summary score (SF-36 PCS), constructed from a combination of responses measured on the 36-Item Short Form Survey (SF-36).
When using a derived variable, a particular complication may arise when some of the measured ‘source’ variables required to derive it have missing values. Missingness in source variables will often lead to missingness in derived variables. For example, if weight is missing for certain study participants, their BMI cannot be derived. However, this is not quite the same as having completely-missing information on the derived variable. In some instances, one might have some information. For example, if a single question on the SF-36 is left unanswered, one might still be able to infer, at least approximately, a participant’s SF-36 PCS based on their answers to the other 35 questions. More clearly, if a composite outcome is defined to be 1 if either component is 1, then observing that one component equals 1 means the composite is known whether or not the other component is measured (Pham et al. 2021).
The data that initially motivated this work consists of measurements taken from newborns of mothers infected with the ZIKA virus (Wilder-Smith et al. 2019). In order to determine if the newborns have microcephaly (a derived variable), a calculation based on three source variables (sex, gestational age, and head circumference) is required. The source variables are often subject to missingness and we were curious as to how one might be able to fit these data with an entirely Bayesian solution.
Besides so-called complete-case analysis, in which observations with any missing values are discarded, the most popular statistical approach for handling missing data is multiple imputation (MI) (Rubin 1976, Reiter & Raghunathan 2007) and among the different varieties of MI, the Multivariate Imputation by Chained Equations (MICE) approach is the most widely used (Van Buuren 2018). The basic idea of MI is to fill in missing values by repeated simulation from the posterior predictive distributions of the missing values given assumptions about the missing data mechanism and a model, thereby preserving the uncertainty associated with the missingness. MICE does this for multivariate non-monotone missing data variables by using a sequence of univariate conditional distributions (i.e., imputing variables one-at-a-time).
When using MICE to impute missing values of derived variables, there are four general approaches:
  1. 1.
    imputation at the derived-variable level (DVL), in which one imputes the derived variable directly ignoring the source variables;
  2. 2.
    imputation at the source-variable level (SVL), in which one imputes the source variables first and then computes the derived variable afterwards (also known as the impute, then transform approach);
  3. 3.
    the just another variable (JAV), approach, in which one imputes the source variables and derived variable together, with the source variables as covariates in the imputation model for the derived variable. To be clear, all variables are imputed separately, ignoring any deterministic relationships. Instead, a multivariate normality between the derived variable and the sources variables is often assumed (e.g., Von Hippel (2009) and White et al. (2011)). Notably, Seaman et al. (2012) conclude that JAV should not be used to impute data for a logistic regression analysis as it performs “very badly” in this case; and
  4. 4.
    the “on-the-fly” imputation, where the transformation of the source variables into the derived variable is done “on-the-fly” within the imputation algorithm. (Note that Van Buuren (2018) calls this “passive imputation”). The “on-the-fly” imputation consists of three steps. First, the source variables are imputed, then the derived variable is updated based on the known deterministic relationship with the source variables, and, third, all other variables are imputed conditional on this updated derived variable. When missingness only occurs in the derived variable and source variables (i.e., no other variables need imputing) and missingness in source variables necessarily implies missingness in the derived variable (e.g., any individual with unknown height or weight will necessarily have unknown BMI), on-the-fly imputation is entirely equivalent to SVL imputation.
The other approach we should mention, proposed by Bartlett et al. (2015), is known as the Substantive Model Compatible Fully Conditional Specification (SCMFCS) imputation. This method is useful for imputing derived covariate values, but not derived outcome variables (which is our focus).
There may be challenges when imputing at the source-variable level if the derived variable is not a linear function of the source variables; see Seaman et al. (2012), Morris et al. (2014), and Tilling et al. (2016). However, because DVL imputation effectively excludes information available in the source variables (which are possibly the best predictors of the missing derived variable values), SVL (or “on-the-fly”) imputation is thought to be a more efficient strategy, in the sense that it uses observed information where possible. Pan et al. (2020) compare the performance of the DVL and SVL approaches and conclude that SVL imputation is indeed preferable, regardless of how the derived variable is obtained (“whether by simple arithmetic operations or by some highly specific algorithms”). This follows related work by Gottschall et al. (2012) who arrived at a similar conclusion for handling missing values in questionnaire data (“researchers should adopt item-level imputation whenever possible”).
$Z_{1}$ (source) $Z_{2}$ (source) $Y$ (derived)
Pattern 1 Missing Observed $\to$ Missing
Pattern 2 Observed Missing $\to$ Missing
Pattern 3 Missing Missing $\to$ Missing
Table 1: Three patterns of missingness in $Z_{1},Z_{2}$ that result in $Y$ being missing, where $Y=Z_{1}+Z_{2}$.
Another reason to favour the SVL, “on-the-fly”, and JAV approaches over DVL imputation is that they make a less restrictive MAR (missing at random) assumption. MAR requires that the conditional probability of missingness depends only on data observed under that pattern (Rubin 1976). Consider for example a derived outcome $Y$ that is equal to the sum of two source variables, $Z_{1}$ and $Z_{2}$, such that $Y=Z_{1}+Z_{2}$. Missingness in $Y$ can occur in three patterns: First, the $Z_{1}$ value is missing, second the $Z_{2}$ value is missing, third, both $Z_{1}$ and $Z_{2}$ values are missing; see Table 1. Because MAR is i) defined at the observation level rather than at the variable level, and ii) conditional on variables included in the analysis, we see that assuming MAR with DVL imputation requires that missingness in the derived variable $Y$ be independent of the value of $Z_{1},Z_{2}$ under any pattern, see Table 2. For the SVL, “on-the-fly”, and JAV approaches, this is relaxed under patterns 1 and 2: missingness in $Z_{1}$ under pattern 1 need only be independent of $Z_{1}$, not $Z_{2}$, and missingness in $Z_{2}$ under pattern 2 need only be independent of $Z_{2}$, not $Z_{1}$.
$A=Z_{1}$ (source) $A=Z_{2}$ (source) $A=Y$ (derived)
The MAR assumption requires that, under this
pattern, missingness in $A$ is independent of
Pattern 1: $Z_{1}$ $Z_{1},Z_{2}$
Pattern 2: - $Z_{2}$ $Z_{1},Z_{2}$
Pattern 3: $Z_{1},Z_{2}$ $Z_{1},Z_{2}$ $Z_{1},Z_{2}$
Table 2: For each of the patterns in table 1, this table shows the what the MAR assumption requires missingness be independent of. For the DVL imputation approach, missingness in the $Y$ must be independent of the value of $Z_{1}$ and the value of $Z_{2}$, regardless of pattern.
To conduct Bayesian inference with imputed data, Gelman et al. (2004) outline a three step procedure. The first step is to construct $K$ imputed datasets, using MI. The second step is to simulate draws from the posterior distribution of the model parameters with each imputed dataset separately. The third step is to mix all the draws together, thereby obtaining (at least approximately) draws from the posterior distribution that take into account the uncertainty associated with the missing data; see Zhou & Reiter (2010). Summaries (e.g., median, 2.5th, and 97.5th percentiles) of the mix of draws from across the imputed datasets are then obtained for one’s parameter estimates. (This third step is perhaps surprising to those familiar with multiple imputation inference: Rubin’s combining rules are not needed.) While Gelman et al. (2004)’s three-step procedure appears straightforward, when SVL (or “on-the-fly”) imputation is used for imputing a derived outcome variable, things can be much more complicated. This becomes apparent when considering the assumptions required in the first two steps.
Using SVL or “on-the-fly” imputation in the first step requires one to specify distributions for each of the missing source variables. Then, specifying the posterior distribution in second step requires one to define a distribution – along with priors for the parameters that characterize the distribution – for the derived outcome variable. Clearly the assumed distributions of the source variables (specified in step 1) necessarily determine the distribution of the derived outcome variable (which must be specified in step 2). Without careful consideration, one could inadvertently define incompatible distributions, since the two steps are done separately and the software implementing one step (e.g., MICE from van Buuren et al. (2015)) does not know of the assumptions required by the other step (e.g., rstan from Carpenter et al. (2017)). Moreover, it will no doubt be difficult to determine whether or not the assumptions made by the Bayesian model specified in the second step (e.g., regarding priors) are consistent with the assumptions made when imputing missing values in the first step.
Ideally, one could fit a single fully Bayesian model to simultaneously impute missing values and sample from the posterior. The idea of avoiding potential problems with MI by adopting a fully Bayesian approach is not new and it is actually quite common to proceed in this way when covariate values are missing; see Bartlett et al. (2015), Lüdtke et al. (2020) and Erler et al. (2021) and the references therein. However, it is not at all obvious how to proceed when there is missingness in source variables that combine to define a derived outcome variable. In this article, we propose a useful approach.
In Section 2, we begin with a simple illustrative example. In Section 3, we outline our proposed approach. In Section 4, we consider two applied examples: (1) comparing the BMI of Dutch boys from inside the city to those from outside the city, and (2) estimating the risk of microcephaly in newborns exposed to the ZIKA virus with an artificial dataset. Finally, we conclude in Section 5.
2 A simple illustrative example
We continue with the simple example discussed in the introduction where $Y=Z_{1}+Z_{2}$. Suppose we are interested in two groups, group $A$ and group $B$, (e.g., placebo and treatment) and we are interested in estimating the difference between the mean of $Y$ in group $A$, and the mean of $Y$ in group $B$. The easiest way to proceed might be to fit a simple univariate Normal model for $Y$ where, for $i=1,\ldots,n$:
$y_{i}$ $\sim\text{Normal}(\alpha+\beta x_{i},\sigma_{y}^{2})\quad$ (1)
where $x_{i}=0$ if observation $i$ is in group $A$, and $x_{i}=1$ if observation $i$ is in group $B$, with priors specified for $\alpha$, $\beta$ and $\sigma_{y}$. The estimand (the target quantity that is to be estimated) coincides with $\beta$.
Now suppose that, in group $A$, a proportion of $Z_{1}$ and $Z_{2}$ values are missing. Specifically, in group $A$, certain observations are prone to have missing $Z_{1}$ values with the chance of missingness depending on $Z_{2}$, and other observations are prone to have missing $Z_{2}$ values with the chance of missingness depending on the observed $Z_{1}$. (To be clear, in this specific example, no observations are susceptible to missing both $Z_{1}$ and $Z_{2}$.) Table 3 lists the data for six individuals to illustrate, with “NA” indicating missing. Given this missingness in the data, how should we proceed to estimate $\beta$?
$i$ X (Group) $Y$ $Z_{1}$ $Z_{2}$
1 A NA NA 1.71
2 A 1.59 --0.08 1.68
3 A NA 1.23 NA
4 B 2.40 2.20 0.20
5 B 4.49 2.35 2.14
6 B 3.53 0.49 3.03
Table 3: Six observations from the hypothetical dataset considered in the example.
As discussed in the introduction, one could consider the complete-case analysis, where all rows of data with any missing values are discarded (individuals $i=1$ and $i=3$ listed in Table 3). The advantage of this strategy is that no imputation step is required and the entire procedure is to simply fit the univariate Normal model. The disadvantage is that this strategy almost certainly involves a loss of efficiency and has the potential for bias.
Alternatively, one could apply Gelman’s three step procedure with one of the three MICE approaches (either the DVL, SVL, or JAV imputation methods) for imputing the missing $Y$ values followed by the univariate Normal model for inference. The advantage of this strategy is that, depending on the specific MICE approach used, one could conceivably obtain unbiased estimates of $\beta$ in an efficient manner. The disadvantage is that this strategy involves two separate steps (an imputation step and an inference step) which may not be consistent with one another.
There is another possible strategy worth considering. Because the math in this scenario is rather simple, one could fit a Bayesian model for the source variables, $Z_{1}$ and $Z_{2}$, and derive the “implied” samples of $\beta$. For instance, we could fit a bivariate Normal model for $Z_{1}$ and $Z_{2}$, where, for $i$ in 1,…,$n$:
$(Z_{1,i},Z_{2,i})$ $\sim\text{Normal}(\boldsymbol{\mu_{i}},\boldsymbol{\Sigma}),$ (2)
where
$\boldsymbol{\mu_{i}}$ =$=\begin{cases}(\mu_{Z1,A},\mu_{Z2,A})&\text{if $i$ is in group $A$}\\ (\mu_{Z1,B},\mu_{Z2,B})&\text{if $i$ is in group $B$}\\ \end{cases}$
$\textrm{and}\quad\quad\boldsymbol{\Sigma}$ $=\begin{pmatrix}\sigma_{Z1}^{2}&\rho\sigma_{Z1}\sigma_{Z2}\\ \rho\sigma_{Z1}\sigma_{Z2}&\sigma_{Z2}^{2}\\ \end{pmatrix},$
and, for each MCMC iteration, compute
$\beta$ $=(\mu_{Z1,A}+\mu_{Z2,A})-(\mu_{Z1,B}+\mu_{Z2,B}).$ (3)
Priors would need to be specified for each of the seven parameters: $\mu_{Z1,A}$, $\mu_{Z2,A}$, $\mu_{Z1,B}$, $\mu_{Z2,B}$ $\sigma_{Z1}$, $\sigma_{Z2}$ and $\rho$.
This strategy might be ideal in the sense that it does not require a separate imputation step (the entire procedure is to simply fit the bivariate Normal model) and does not disregard a large number of observations. However, it is possible only because in this simple illustrative example the math is extremely straightforward (i.e., equation (3) is known). For other derived variables, this might not be the case. For instance, if instead of $Y=Z_{1}+Z_{2}$, we had $Y=Z_{1}/({Z_{2}}^{2})$, then computing $\beta$ from a combination of the seven model parameters ($\mu_{Z1,A}$, $\mu_{Z2,A}$, $\mu_{Z1,B}$, $\mu_{Z2,B}$ $\sigma_{Z1}$, $\sigma_{Z2}$ and $\rho$) might be impossible without a degree of mathematical wizardry.
This brings us to our proposed method, details of which are provided in the next section. Briefly, the approach involves using Monte Carlo integration to derive samples of the derived outcome variable. In our example this would involve fitting the bivariate Normal model and sampling $\beta$ (at least approximately) without requiring any knowledge about the mathematical derivation of $\beta$.
3 Description of the approach
Having established the motivation behind our proposed method, we now outline the details.
Suppose the data consists of $n$ observations and let $W$ be the set of $p+q$ variables at hand. Let $Y=f(W)=f(W_{source})$ be the outcome variable, where $W_{source}$ are the $p$ source variables that combine in some deterministic way to equal the outcome variable. (To be clear, there are $p$ source variables which make up the outcome $Y$). Let $W_{expl}$ consist of the remaining $q$ explanatory variables. In practice, note that certain variables could be both source variables and explanatory, but we set this aside to avoid more complicated notation. Here, $\dim(W_{source})=(n,p)$ and $\dim(W_{expl})=(n,q)$.
Say $W$ partitions as variables $W_{1}$ which are prone to missingness, and variables $W_{2}$ which are not: $W=\{W_{1},W_{2}\}$. Specifically, $W_{1}$ consists of certain observations that are missing, $W_{1}^{\{mis\}}$, and certain observations that are observed, $W_{1}^{\{obs\}}$, such that $W_{1}=\{W_{1}^{\{mis\}},W_{1}^{\{obs\}}\}$. We are interested in the situation in which there is overlap between $W_{1}$ and $W_{source}$, i.e., situations when certain source variables are subject to missingness. Let $W_{1,source}=W_{1}\cap W_{source}$, $W_{1,expl}=W_{1}\cap W_{expl}$, $W_{2,source}=W_{2}\cap W_{source}$, and $W_{2,expl}=W_{2}\cap W_{expl}$.
In the simple illustrative example we have $W=\{Z_{1},Z_{2},X\}$ (where $X$ is the binary group indicator variable such that, for the $i$-th observation, $X_{i}=0$ if $i$-th observation is in Group A and $X_{i}=1$ if the $i$-th observation is in Group B). Furthermore, we have $W_{1}=\{Z_{1},Z_{2}\}$, $W_{2}=\{X\}$, $p=2$, $W_{source}=\{Z_{1},Z_{2}\}$, $q=1$, $W_{expl}=\{X\}$, and $Y=f(W)=Z_{1}+Z_{2}$.
Typically, the objective of a Bayesian analysis consists of the estimation of an estimand (Kahan et al. 2023). The estimand of interest, $\theta$, is often defined non-parametrically as a function of certain conditional expectations: $\theta=g(\textrm{E}^{*}_{1}\{\textrm{E}(Y|W_{expl})\},...,\textrm{E}^{*}_{D}\{\textrm{E}(Y|W_{expl})\})$, where $\textrm{E}^{*}_{d}$ corresponds to the counterfactual expectation with respect to the $d$-th population of interest for $d$ in 1,…,$D$.
In certain situations, the target population(s) might themselves be unknown and require estimation. For example, suppose the estimand of interest is the mean difference in $Y$ between those receiving treatment ($X=1$) and those not receiving treatment ($X=0$) for individuals who are of “average age”, where average age, $\mu_{A}$, is unknown. Then, we have $D=2$ target populations: (1) treated individuals of average age, and (2) untreated individuals of average age. The estimand is defined as: $\theta=\textrm{E}(Y|X=1,A=\mu_{A})-\textrm{E}(Y|X=0,A=\mu_{A})$. The parameter $\mu_{A}$ is required to define the two target populations and must be estimated from the data.
In the simple illustrative example, the estimand of interest is the difference in the mean of $Y$ between individuals in the two groups (i.e., the difference between the conditional expectation of $Y$ given the $D=2$ target populations defined by $X=0$ and $X=1$): $\theta=\textrm{E}(Y|X=1)-\textrm{E}(Y|X=0)$. Because the math is conveniently simple, one can also define this estimand analytically as a function of the model parameters: $\theta=\beta$, but also as: $\theta=h(\phi_{source})=(\mu_{Z1,A}+\mu_{Z2,A})-(\mu_{Z1,B}+\mu_{Z2,B})$, where $\phi_{source}=\{\mu_{Z1,A}$, $\mu_{Z2,A}$, $\mu_{Z1,B}$, $\mu_{Z2,B}$ $\sigma_{Z1}$, $\sigma_{Z2}$, $\rho\}$. The standard Bayesian approach would then be to obtain a Monte Carlo sample of size $M$ (e.g., via MCMC) from the $(\phi_{source},W_{1}^{\{mis\}}|W_{1}^{\{obs\}},W_{2})$ posterior distribution and then, since $h()$ is known, obtain $M$ posterior samples of $\theta$ by applying the $h()$ function to each of the $M$ posterior samples of $\phi_{source}$.
In situations when the math is not so simple, Bayesian g-computation (Gustafson 2015, Keil et al. 2018) can be used for a form of ‘model-based standardization’ (Zou 2009, Vansteelandt & Keiding 2011) in order to obtain posterior samples of $\theta$ as follows. First, for each of the target populations of interest, one forward samples “posterior predictive values” of the source variables conditional on the target population of interest. Then, by applying the deterministic function $f()$ on the sample of these “posterior predictive values” one can obtain a sample of $Y^{*}_{d}$, a new variable with the same identical distribution as $Y$ under the $d$-th population of interest, i.e., a “future observation” projected by the model of the $Y$ variable from the $d$-th target population. Finally, with a large number, $S$, of these “future observations” one can approximate $\textrm{E}^{*}_{d}(\textrm{E}(Y|W_{expl}))$, for $d$ in 1,…$D$, and combine these according to the definition of the estimand of interest to obtain a posterior sample of $\theta$. Obtaining samples of $\theta$ in this way ensures that the correct amount of variation is captured and, importantly, does not require knowledge of a function $h:\phi_{source}\rightarrow\theta$.
To be clear this approach consists of the following steps:
  1. 1.
    Draw $M$ samples of $(\phi_{source},W_{1,expl}^{\{mis\}})$ from the $(\phi_{source},W_{1,expl}^{\{mis\}}|W_{1}^{\{obs\}},W_{2})$ distribution: $(\phi_{source},W_{1,expl}^{\{mis\}})^{[m]}$, for $m$ in 1,…,$M$.
  2. 2.
    For $m$ in 1,…,$M$:
    1. (a)
      For $d$ in 1,…,$D$:
      1. i.
        If the $d$-th target population is defined in terms of unknown parameters, $\gamma^{d}$, these must be estimated from the data. Obtain a draw from $\gamma^{d}|(\phi_{source},W_{1,expl}^{\{mis\}})^{[m]}$: $\gamma^{d[m]}$.
      2. ii.
        Draw $S$ (where $S$ is a large number) iid samples from the $(W_{source}|\gamma^{d[m]})$: $W^{*d[m,s]}_{source}$, for $s$ in 1,…, $S$. Note: $\dim(W^{*d[m,s]}_{source})=(1,P)$.
      3. iii.
        For $s$ in 1,…,$S$: Apply the deterministic function $f()$ to $W^{*d[m,s]}_{source}$ to obtain a sample from the $(Y|\gamma^{d[m]})$ distribution: $Y^{*d[m,s]}=f(W^{*d[m,s]}_{source})$. Note: $\dim(Y^{*d[m,s]})=(1,1)$.
      4. iv.
        Calculating (approximately) the sample mean:
        $\textrm{E}^{*}_{d}\{\textrm{E}(Y|W_{expl})\}^{[m]}\approx\frac{1}{S}\sum_{s=1}^{S}{Y^{*d[m,s]}}$.
    2. (b)
      Obtain ${\theta}^{[m]}$ by combining $\textrm{E}^{*}_{1}\{\textrm{E}(Y|W_{expl})\},...,\textrm{E}^{*}_{D}\{\textrm{E}(Y|W_{expl})\}$ according to the definition of the estimand of interest.
Having obtained $M$ posterior samples of $\theta$ following the four steps, one can calculate the posterior median/mean and credible interval for $\theta$, as is standard practice in a Bayesian analysis. Note that both $M$ and $S$ should be sufficiently large that both sources of Monte Carlo error are negligible.
In the simple illustrative example, the $D=2$ target populations were defined by $X=0$ and $X=1$. Therefore the $m$-th posterior sample of $\theta$ is approximated as:
$\theta^{[m]}=\frac{1}{S}\sum_{s=1}^{S}{Y^{*2[m,s]}}-\frac{1}{S}\sum_{s=1}^{S}{Y^{*1[m,s]}},$ (4)
where $Y^{*1[m,s]}$ is a sample from the distribution of $(Y|\phi_{source}^{[m]},X=0)$ and $Y^{*2[m,s]}$ is a sample from the distribution of $(Y|\phi_{source}^{[m]},X=1)$.
4 Applied examples
4.1 Multiple linear regression of the body mass index of Dutch boys
Background
The “Growth of Dutch boys” dataset is available as an example dataset within the ‘mice’ package for R (van Buuren et al. 2015) and is often used to demonstrate how various imputation methods can be applied. Here we use this dataset to illustrate how the approach outlined in Section 3 could be applied for determining the difference in logBMI between boys from the “city” and boys from outside of the “city”, when controlling for age.
The outcome of interest, $Y=\textrm{log}(\textrm{BMI})$, is a derived binary variable, based on two source variables: (1) height ($Z_{1}$, measured in centimeters), (2) weight ($Z_{2}$, measured in kilograms). This is derived as follows:
$Y=\textrm{logBMI}=f(Z_{1},Z_{2})=\textrm{log}\Big{(}\frac{Z_{2}}{({Z_{1}/100})^{2}}\Big{)}.$ (5)
The exposure variable is regional status ($X$, equal to 1, if the boy is from the “city” and equal to 0 otherwise). We also consider the variable age ($A$, measured in years) as highly predictive of $Z_{1}$, $Z_{2}$, and $Y$.
The data
The data consists of measurements on a representative sample of $n=537$ boys between the ages of 1 and 18 years (inclusive), of which approximately 9% are from the “city” (we restricted the original “Growth of Dutch boys” dataset dataset ($n=748$) to the 1–18 age range to allow for reasonable modeling of how log(BMI) changes with age). Figure 1 plots the data among the complete cases. All variables, except age ($A$), are subject to a small amount of missingness. The exposure variable, “city” ($X$), is missing for 1 subject; height is missing for 18 subjects; and weight is missing for 2 subjects; see Venn diagram in Figure 2. Overall, 19 out of the 537 of observations (3.5%) have at least one missing value. Following the notation defined in Section 3, we have that $W_{1}=\{Z_{1},Z_{2},X\}$, $W_{2}=\{A\}$, $p=2$, $W_{source}=\{Z_{1},Z_{2}\}$, $q=2$, and $W_{expl}=\{X,A\}$.
The primary objective of the study is to estimate the difference in mean logBMI between boys from inside the “city” and boys from outside of the “city”, when controlling for age. A secondary objective is to make probabilistic predictions of logBMI for boys given their age and regional status.
Figure 1: The complete case data for the Dutch boys analysis ($n$=518) with violet dots corresponding to boys from outside the city ($X=0$) and green dots corresponding to boys from the city ($X=1)$. Lines drawn from least-squares estimates.
Figure 2: Venn diagram for missingness in Dutch boys dataset ($n$=537).
The estimand
Based on the primary objective, the estimand of interest is the difference in mean logBMI between boys from the “city” and boys from outside of the “city”, when controlling for age, and this can be defined as a function of the conditional expectation of $Y$ given $D=2$ target populations:
[…]
where $f_{A}(a)$ corresponds to the distribution of ages among boys in the Netherlands. Since the data supposedly consist of a representative sample of Dutch boys, we assume that $f_{A}(a)$ also coincides with the sample distribution of ages.
Two models
The univariate normal model involves considering the logBMI of each subject as the outcome of a simple univariate Normal model:
$(Y|X=x,A=a)\sim\textrm{Normal}(\beta_{0}+\beta_{1}x+\beta_{2}a+\beta_{3}xa+\beta_{4}a^{2},\sigma^{2}).$ (7)
This model assumes a quadratic relationship between BMI and age, a reasonable approximation based on published growth curves (Kuczmarski 2000) and allows for the possibility that age is an effect modifier with respect to the effect of “city”. For the univariate normal model, we set standard Normal priors, $\textrm{Normal}(0,1)$, for the regression coefficient parameters $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$ and $\beta_{4}$, an exponential prior, $\textrm{Exponential}(1)$, for the standard deviation parameter, $\sigma$, and a uniform prior, $\textrm{Uniform}[0,1]$, for $\pi$.
The bivariate normal model considers height and weight as two correlated outcomes such that:
$(\textrm{log}(Z_{1}),\textrm{log}(Z_{2})|X=x,A=a)$ $\sim\text{Normal}(\boldsymbol{\mu_{i}},\boldsymbol{\Sigma}),$ (8)
where
$\boldsymbol{\mu_{i}}$ $=\begin{pmatrix}\alpha_{0}+\alpha_{1}r+\alpha_{2}a+\alpha_{3}xa+\alpha_{4}a^{2}\\ \gamma_{0}+\gamma_{1}x+\gamma_{2}a+\gamma_{3}xa+\gamma_{4}a^{2}\end{pmatrix}\textrm{and}\quad\quad$ (9)
$\boldsymbol{\Sigma}$ $=\begin{pmatrix}\tau_{Z1}^{2}&\rho\tau_{Z1}\tau_{Z2}\\ \rho\tau_{Z1}\tau_{Z2}&\tau_{Z2}^{2}\\ \end{pmatrix}.$
and:
$R\sim\textrm{Bernoulli}(\pi).$ (10)
We set standard Normal priors, $\textrm{Normal}(0,1)$, for all the regression coefficient parameters $\alpha_{0}$, $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, $\alpha_{4}$, $\gamma_{0}$, $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$; exponential priors, $\textrm{Exponential}(1)$, for the standard deviation parameters, $\tau_{Z1}$ and $\tau_{Z2}$, a uniform prior, $\textrm{Uniform}[-1,1]$, for the correlation parameter, $\rho$, and another uniform prior, $\textrm{Uniform}[0,1]$, for $\pi$.
We can define the estimand of interest analytically as a function of the parameters of the univariate normal model:
[…]
$=\beta_{1}+\beta_{3}\textrm{E}(A),$ (12)
where $\phi_{univ}=\{\beta_{0},\beta_{1},\beta_{2},\beta_{3},,\beta_{4},\sigma\}$. Since the data is assumed to be a representative sample from the target population, and since we are choosing not to model the distribution of age (not required because there are no missing values for age), we can substitute $\textrm{E}(A)$ for the sample mean, $\bar{A}$.
Coincidentally, we can define the estimand as a function of the parameters of the bivariate normal model (see details in the Appendix):
[…]
$=\gamma_{1}-2\alpha_{1}+(\gamma_{3}-2\alpha_{3})\textrm{E}(A),$ (13)
where $\phi_{biv}=\{\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\tau_{Z1},\tau_{Z2},\rho\}$, and we once again can substitute $\textrm{E}(A)=\bar{A}$.
Methods and Results
We conducted the analysis with four different approaches:
  1. 1.
    Fitting the univariate normal model to the complete case data and applying the $h_{1}()$ function to obtain posterior draws of $\theta$ (with $M=2,000$).
  2. 2.
    Fitting the univariate normal model with Gelman’s three-step approach with “on-the-fly” multiple imputation (including a quadratic term for age and age-city interaction term in the imputation model), and applying the $h_{1}()$ function to obtain posterior draws of $\theta$ (with $M=2,000$, and $K=50$ imputed datasets).
  3. 3.
    Fitting the bivariate normal model and applying $h_{2}()$ function to obtain posterior draws of $\theta$ (with $M=2,000$).
  4. 4.
    Fitting the bivariate normal model and applying the proposed method detailed in Section 3 (with $M=2,000$ , and $S=2,000$).
Both the univariate model and bivariate model were fit using JAGS with 1,000 burn-ins (Plummer 2004). R and JAGS code is available in the Appendix.
Results are listed in Figure 3. The estimate from the complete case analysis is similar to the estimates obtained when using the the bivariate model with a wider credible interval, suggesting that one can gain precision by incorporating the small number of observations with missing values into the analysis. However, one obtains a wider credible interval when using the univariate normal model with Gelman’s three-step approach with “on-the-fly” multiple imputation, suggesting that precision can also be lost. Finally, the estimate obtained with fitting the bivariate normal model and applying $h_{2}()$ function is slightly more precise than the estimate obtained with fitting the bivariate normal model and applying the proposed method, suggesting that there is a small cost to the numerical approximation (even with the large $S=2,000$). With respect to computational time, the proposed method is substantially more time consuming than the other methods.
Figure 3: Results for the analyses of the Dutch boys dataset. Estimates list the posterior median estimates (i.e., the median of the posterior estimates of the mean difference in logBMI) and equal-tailed posterior 95%CrIs. Labels indicate computational time required for each method in minutes.
4.2 Estimating the risk of microcephaly
Background
Zika virus infection (ZIKV) during pregnancy is known to be associated with an increased risk of fetal congenital malformations, including microcephaly, a birth condition in which a baby’s head is smaller than expected when compared to babies of the same sex and age. Formally, microcephaly is defined as a measure of head circumference at birth of more than 2 standard deviations below average for sex and gestational age; see Harville et al. (2020). (Average values for head circumference at birth are determined by referencing the standards calculated by the intergrowth-21st research network; see Villar et al. (2014)). Therefore, in a healthy population, we would anticipate 2.28% of newborns being categorized as having microcephaly (assuming head circumference in the general population is Gaussian).
In order to determine the risk of microcephaly associated with ZIKV in mothers, observational data has been collected by numerous studies from infected pregnant women (Zika Virus Individual Participant Data Consortium 2020). These data typically suffer from substantial missingness.
The outcome of interest, microcephaly ($Y$), is a derived binary variable, based on three source variables: (1) sex ($Z_{1}$=0 for ‘male’ and $Z_{1}$=1 for ‘female’), (2) gestational age ($Z_{2}$, measured in weeks), and (3) head circumference ($Z_{3}$, measured in centimetres). Each of these three source variables is subject to missingness. The estimand of interest is the probability of microcephaly ($\theta=\textrm{E}(Y)$) given then assumed $D=1$ target population from which the data is sampled.
The data and models
We consider a simple synthetic dataset based on what is typically observed in observational studies. The purpose of this analysis is to demonstrate all of the steps required when following each of the different approaches for dealing with missingness based on the problem that initially motivated our research. As such, we stress that the specific details of how the data were created are less important and acknowledge that a different dataset would no doubt lead to different results. In brief, the data is simulated with a “true” risk of microcephaly of $\theta$ = 11.69%. The data represent $Z_{1},Z_{2},Z_{3}$ measurements on $n=1,800$ newborns with at least one of $Z_{1},Z_{2},Z_{3}$ observed; Figure 4 shows $Z_{3}$ and $Z_{2}$ among those with both observed. Missing $Z_{1}$ values are attributed completely at random, missing $Z_{2}$ values are more likely for those with very small values of $Z_{3}$, and missing $Z_{3}$ values are more likely for those with very large values of $Z_{2}$.
There are 781 individuals who have at least one value missing; see Venn diagram in Figure 5. Microcephaly status, $Y$, for a newborn can be calculated by first deriving their “z-score” based on their head circumference, gestational age and sex. (The “igb_hcircm2zscore” function in the growthstandards R package can do this easily using the intergrowth standards). If a newborn’s z-score is below $-2$, they are classified as having microcephaly. As such, while the $f()$ function is known and can be easily applied to the source variables to obtain the binary derived variable, it is not easily written out.
We consider two models: (1) a simple “Bernoulli model”, and (2) a “Bernoulli skew-Normal mixture-Normal model” (BsNmN model). Both models were fit using Stan with $M=5,000$ draws (6,000 iterations and 1,000 burn-ins) (Carpenter et al. 2017); Stan code is available in the Appendix.
Figure 4: The complete case data ($n$=1,019) with values (1) sex ($Z_{1}$=0 for ‘male’ and $Z_{1}$=1 for ‘female’), (2) gestational age ($Z_{2}$, measured in weeks), and (3) head circumference ($Z_{3}$, measured in centimetres).
Figure 5: Venn diagram and UpSet plot indicating the number of observations in the dataset with observed data. There are $n$=1,019 individuals for which there is complete data and $n$=781 individuals who have at least one value missing.
The Bernoulli model involves considering the binary microcephaly status of each individual as the outcome of a simple Bernoulli model where $\theta$ is the risk of microcephaly:
$Y\sim\textrm{Bernoulli}(\theta).$ (14)
We set a $\textrm{Beta}(1.26,2.32)$ prior for $\theta$ (see top panel of Figure 6), corresponding to an a priori belief that $\textrm{Pr}(\theta<31.5\%)=50\%$. (This prior for was selected so that the two models have (implied) priors matching as closely as possible; see Figure 6).
We fit the Bernoulli model to the complete case data ($n=1,019$) and obtained an estimate of: $\hat{\theta}=9.18\%$, 95%CrI= [7.60%, 10.92%] (posterior median and equal-tailed 95% credible interval). We also fit the Bernoulli model to the entire dataset ($n$= 1,800) and followed Gelman’s three-step procedure with SVL imputation. Specifically, we used MICE specifying the “logreg” method for $Z_{1}$ and the “normal” method for $Z_{2}$ and $Z_{3}$. For each of $K=50$ imputed datasets, we fit the Bernoulli model using MCMC with $M$ = 5,000 draws and after combining all 250,000 draws together, we obtained a posterior median estimate of: $\hat{\theta}=12.70\%$, with equal-tailed 95%CrI= [11.00%, 14.51%].
The BsNmN model involves a series of conditional distributions for each of the three source variables. We assume that an infant’s sex ($Z_{1}$) is unlikely to impact their gestational age at birth ($Z_{2}$) (at least by any appreciable amount; see Broere-Brown et al. (2016)). Therefore, for sex, we define:
$Z_{1}\sim\textrm{Bernoulli}(0.5).$ (15)
The distribution of gestational age ($Z_{2}$) is known to be skewed with a long left tail (due to preterm delivery), almost complete truncation on the right tail at 45 weeks (due to medically-induced labour at around 45 weeks) and almost complete truncation on the left tail at around 24 weeks (due to viability). Many complex distributions have been suggested for modelling gestational age (e.g., Rathjens et al. (2023) recommends the three-parameter Dagum distribution). Following Sauzet et al. (2015), we choose to define a skew-Normal distribution centered at 39 weeks:
$(Z_{2}-39)\sim\textrm{skew-Normal}\left(\mu-\sqrt{\frac{2}{\pi}}\frac{\sigma\omega}{\sqrt{1+\omega^{2}}},\sigma,\omega\right),$ (16)
where $\mu$ is the mean of the distribution (known to be approximately 39 weeks, according to Sauzet et al. (2015)), $\sigma$ is the scale, and $\omega$ is the slant.
Finally, head circumference ($Z_{3}$) is known to be approximately Gaussian and mean head circumference increases in an approximately quadratic way with gestational age (see Figure 2C in Villar et al. (2014), and Figure 11 in the Appendix). In a ZIKV infected population, infection is thought to cause a certain proportion of infants to have smaller heads. As in Kalmin et al. (2019), we define a Normal mixture distribution for $Z_{3}$:
$Z_{3}|Z_{1}=z_{1},Z_{2}=z_{2},X=1\quad$ $\sim$
$\quad w\textrm{Normal}(\beta_{0}+\beta_{1}z_{1}$ $+\beta_{2}(z_{2}-39)+\beta_{3}(z_{2}-39)^{2},\zeta_{1})+$
$(1-w)\textrm{Normal}(\kappa$ $+\beta_{1}z_{1}+\beta_{2}(z_{2}-39)+\beta_{3}(z_{2}-39)^{2},\zeta_{2}),$ (17)
where $w$ is the proportion of “non-affected” individuals and $(1-w)$ is the proportion of “affected” individuals. The $\beta_{0}$, $\beta_{1}$, $\beta_{2}$ and $\zeta_{1}$ parameters relate the distribution of head circumference to sex and gestational age for the “non-affected” population; and the $\kappa$ and $\zeta_{2}$ parameters correspond to the impact of ZIKV for the “affected” individuals in terms of the difference in mean head circumference and the scale.
The parameters in the BsNmN model are $\psi=(\mu,\sigma,\omega,\beta_{0},\beta_{1},\beta_{2},\beta_{3},\zeta_{1},\zeta_{2},\kappa,w)$ and each requires a prior. Based on the information available in Villar et al. (2014) and Sauzet et al. (2015), we can be reasonably certain of how gestational age and head circumference are distributed in the non-affected population and therefore set the following informative priors:
$\mu\sim\textrm{Normal}(0,0.1),$
$\beta_{0}\sim\textrm{Normal}(33.912,0.1),\quad$ $\beta_{1}\sim\textrm{Normal}(-0.450,0.1),$
$\beta_{2}\sim\textrm{Normal}(0.399,0.1),\quad\textrm{and}$ $\beta_{3}\sim\textrm{Normal}(-0.016,0.1).$
For the scale and slant parameters, we set the following weakly informative priors:
[…]
Finally, a truncated Normal prior for $\kappa$ was chosen in an effort to help address issues with identifiability that are common when fitting Bayesian mixture models (Betancourt 2017), and a uniform prior was chosen for $w$:
$\kappa$ $\sim\textrm{Normal}(-2,2)_{[,-1]},\quad\quad w\sim\textrm{Uniform}(0,1).$
These priors correspond to an a priori belief that $\textrm{Pr}(\textrm{E}(Y)<31.5\%)=0.5$. The bottom panel of Figure 6 plots the implied prior distribution for $\textrm{E}(Y)$ (i.e., for the risk of microcephaly) and Figure 7 plots the implied distribution of the z-score from 12 random draws of the prior.
Fitting the BsNmN model to the complete case data ($n=1,019$) we obtain an estimate of: $\hat{\theta}=9.18\%$, 95%CrI= [7.60%, 10.92%]. We also applied the proposed approach to fit the BsNmN model to the entire dataset ($n=1,800$) as detailed in Section 2 with $S$=5,000 and obtained an estimate of: $\hat{\theta}=11.68\%$, 95%CrI= [10.16%, 13.36%]. Figure 9 plots the implied distribution of the z-score from the prior and posterior. Figures 12 and 13 in the Appendix show the trace plots of the MCMC.
Figure 6: Lower panel shows histogram of a sample from the “implied prior” on $\textrm{E}(Y)$ used in the BsNmN model for the microcephaly example analysis. Upper panel shows histogram of a sample from the Beta(1.26, 2.32) prior placed on $\theta$ used in the Bernoulli model. The numbers 1.26 and 2.32 were specifically chosen so that these two priors would be similar.
Figure 7: The implied distribution of the z-score from 12 random draws from the prior of the BsNmN model for the microcephaly example analysis. For reference, the red curve corresponds to the density of the standard normal.
Figure 8: Results from the microcephaly analyses (posterior medians and equal-tailed 95% credible intervals).
Figure 9: The implied distribution of the z-score from 12 random draws from the posterior of the BsNmN model. Newborns with z-scores less than $-2$ (vertical dashed line) are classified as having microcephaly. For reference, the red curve corresponds to the density of the standard normal and the blue curve corresponds to the “truth” from which the data were simulated.
5 Discussion
If the amount of missingness is very small or negligible, applying a missing value procedure should not drastically change one’s estimates, even if this procedure involves a misspecified model for the missing values. Within Gelman’s three step method, if the amount of missingness is negligible, an incorrect imputation model should have only a minimal negative effect. In our proposed fully-integrated Bayesian approach, the model for the missing values and the outcome model are one and the same. Therefore if this model is misspecified, the consequences might be more substantial even if there is negligible missingness. This model might also be more complex than the one originally considered for analysis. For instance, a researcher might be able to make relatively simple assumptions about the distribution of the derived outcome variable (e.g., be able to assume that microcephaly status is binomial), but have difficulty in making reasonable distributional assumptions for the source variables (e.g., what are reasonable distributions to assume for gestational age and head circumference?).
The advantages of modeling the source variables directly include the ability to incorporate informed priors specific to the individual source variables and model missingness within a fully Bayesian model following our proposed method. One could also incorporate adjustments for other sources of bias beyond missingness such as preferential sampling (Campbell et al. 2022) and measurement error. This is an important future goal for the microcephaly model given the issues considered by Harville et al. (2020).
Finally, we note that our proposed approach allows one to conduct the analysis of data with missingness within an entirely Bayesian framework. In some instances, one might wish to fit a Bayesian imputation model for the missing values and frequentist model for the main analysis (i.e., for the outcome model). In such a case, Rubin’s rules could be appropriately and easily applied. Alternatively, if both the imputation model and the outcome model of interest must be frequentist, von Hippel & Bartlett’s rules (Von Hippel & Bartlett 2021) may be more appropriate. To the best of our knowledge, if the imputation model is frequentist and the outcome model is Bayesian (an admittedly obscure scenario), there does not appear to be a good approach for obtaining estimates.